# Stop Wasting Your Tokens: Towards Efficient Runtime Multi-Agent Systems

### v1

@@ Abstract
While Multi-Agent Systems (MAS) excel at complex tasks, their growing autonomy with operational complexity often leads to critical inefficiencies, such as excessive token consumption and failures arising from misinformation. Existing methods primarily focus on post-hoc failure attribution, lacking proactive, real-time interventions to enhance robustness and efficiency. To this end, we introduce SupervisorAgent, a lightweight and modular framework for runtime, adaptive supervision that operates without altering the base agent’s architecture. Triggered by an LLM-free adaptive filter, SupervisorAgent intervenes at critical junctures to proactively correct errors, guide inefficient behaviors, and purify observations. On the challenging GAIA benchmark, SupervisorAgent reduces the token consumption of the Smolagent framework by an average of 29.45% without compromising its success rate. Extensive experiments across five additional benchmarks (math reasoning, code generation, and question answering) and various SoTA foundation models validate the broad applicability and robustness of our approach. The code is available at https://github.com/LINs-lab/SupervisorAgent.
@@ 1 Introduction
The advent of powerful Large Language Models (LLMs) has catalyzed significant advancements in Multi-Agent Systems (MAS) (Liu et al., 2025a; ang Gao et al., 2025), enabling them to achieve remarkable performance across diverse and challenging domains such as mathematical reasoning (Shang et al., 2025), code generation (Lu et al., 2025), and complex question answering (Luo et al., 2025). This progress has spurred research into sophisticated agent architectures, including self-evolving systems that learn from feedback and experience (Shi et al., 2025b; Liu et al., 2025b), and dynamic topologies that adapt to task complexity (Li et al., 2025a; b). However, a critical paradox has emerged: as these systems grow more capable and complex, they often become less robust and economically viable (Wu et al., 2025a; Huang et al., 2025). Systemic inefficiencies incur prohibitive computational costs, while intricate interactions introduce vectors for unpredictable failures (Zhang et al., 2025e).
@@
Furthermore, the issue of economic inefficiency is a major barrier to the real-world deployment of MAS (Wang et al., 2025a). We identify two primary sources of this inefficiency. First, agents often struggle with excessively long observations, such as verbose web pages or tool outputs, which flood their context windows. This not only inflates token costs but can also obscure critical information, causing the agent to lose focus and derail its task execution (Hosseini et al., 2025). Second, agents may adopt sub-optimal strategies, entering into repetitive action loops or choosing unnecessarily complex paths to a solution (Cemri et al., 2025), further wasting computational resources (see Figure 1a). Our own analysis reveals that cutting-edge architectures can incur staggering token costs, with some GAIA benchmark (Mialon et al., 2023) tasks consuming upwards of two million tokens.
@@
    We validate the general applicability of our approach across five additional benchmarks spanning mathematical reasoning (GSM8k-Hard, AIME), code generation (HumanEval, MBPP), and question answering (DROP). Our method consistently delivers substantial efficiency gains, highlighted by a 23.74% token reduction on HumanEval alongside an accuracy improvement. The framework’s effectiveness is further confirmed across various foundation models, including the GPT-4.1, Gemini-2.5-pro, and Qwen3 series.
@@ 4 Methodology
Building upon the formalism of a Supervised Multi-Agent System (SMAS) introduced in Section 3, we now detail the architecture and operational workflow of our SupervisorAgent (illustrated in Figure 2). Our methodology is structured around three fundamental questions: What to supervise, When to supervise, and How to supervise. We defer the specific implementation details, including all hyperparameters and prompts, to Appendix A.3 and A.5.
@@ 5.1 Experimental Setup
On several benchmarks, we compare SupervisorAgent against a comprehensive set of agentic systems equipped with web-browsing and code execution capabilities. These baselines fall into two categories: (1) single agent execution methods: including vanilla LLM, Self Consistency CoT (3 answers) (Wang et al., 2023), and CodeAgent (Roucher et al., 2025); and (2) multi-agent systems, including Smolagent (Roucher et al., 2025), OAgents (Zhu et al., 2025), MetaAgent (Zhang et al., 2025h), OWL (role playing) (Hu et al., 2025), and AWorld (Xie et al., 2025; Yu et al., 2025).
@@
To assess model-agnosticism, we test SupervisorAgent with multiple foundation models. For the demanding GAIA benchmark, we primarily use GPT-4.1 as the base model for all agents (see Table 1), and evaluate SupervisorAgent when powered by GPT-4.1 (OpenAI, 2025), Gemini-2.5-pro-0605 (Team, 2025a), and Qwen3-235B-2507 (Team, 2025b); see Figure 4(b). For all other benchmarks (see Table 2), we employ the efficient and powerful Qwen3-32B (Team, 2025b) for both the base agents and the SupervisorAgent to assess performance in a more resource-constrained setting.
@@
Table 1: Overall performance on the GAIA validation set (based on GPT-4.1). Our SMAS consistently reduces the average token cost comparing to Smolagent baseline while achieving competitive success rates.
@@ 5.2 Results and Analysis
Table 2: Generalization across diverse benchmarks (based on Qwen3-32B). SupervisorAgent consistently reduces token costs while maintaining or improving accuracy on tasks spanning mathematical reasoning, code generation, and question answering. All reported gains are relative to the Smolagent baseline.
@@
Our ablation study, conducted on the subset of GAIA benchmark (top-10 most token-intensive tasks per GAIA level), dissects the contribution of SupervisorAgent’s three core strategies (Table 3, Figure 4(a)). We compare the full SupervisorAgent against three ablated versions: w/o Correction (disabling Proactive Error Correction), w/o Guidance (disabling Guidance for Inefficiency), and w/o Purification (disabling Adaptive Observation Purification). The results reveal that Purification is the primary driver of token reduction. Concurrently, removing the Correction and Guidance modules leads to a significant drop in accuracy, confirming their critical role in ensuring task success and robustness. This highlights a clear trade-off: while Purification is key to efficiency, the other two strategies are crucial for maximizing performance, justifying their marginal token cost by preventing costly failures.
Table 3: Ablation study of SupervisorAgent’s components (based on GPT-4.1) on the subset of GAIA benchmark (top-10 most token-intensive tasks per GAIA level), dissecting the contributions of each module.
@@
Smolagent 40.00 1,446,526 933,013 2,037,437 1,369,131
@@
To verify MAS-agnostic feature of SupervisorAgent, we integrate SupervisorAgent into two distinct multi-agent system frameworks: AWorld (Xie et al., 2025) and OAgents (Zhu et al., 2025), and evaluate their performance on the subset of GAIA benchmark(the same as in ablation study). The results, presented in Table 4, indicate that SupervisorAgent consistently enhances the performance of both frameworks, underscoring its versatility and effectiveness across different MAS architectures.
Specifically, integrated with AWorld (Xie et al., 2025), our SMAS(AWorld) achieved superior average accuracy over both the original AWorld (without Guard) and the Guard-enabled version. Furthermore, SMAS(AWorld) demonstrated substantial token efficiency, saving 36.54% on average versus AWorld (with Guard). With savings reaching 48.38% on Level 3 tasks, it confirms SupervisorAgent’s ability to enhance sophisticated MAS.
Applying SupervisorAgent to OAgents (Zhu et al., 2025) further validated its general applicability, reducing average token consumption by 39.36% while maintaining competitive accuracy. Interestingly, the largest token reduction (50.19%) occurred on Level 1 tasks, exceeding the savings on Level 3 (40.63%). While this might reflect OAgents’ inherent proficiency on harder tasks, the significant overall savings underscore SupervisorAgent’s broad utility.
Table 4: Cross-framework performance of SupervisorAgent on Smolagent, AWorld, and OAgents (based on GPT-4.1). Evaluated on the same GAIA subset (top-10 token-intensive tasks per level) as in Table 3.
@@ 6 Discussion
These insights inform several promising avenues for future work. While our method proved efficient and robust, developing a self-evolving, memory-augmented version remains an exciting opportunity. Other key directions include creating more sophisticated purification techniques that address the “noise-as-signal” problem and developing a universal resource consumption metric for MAS. Ultimately, we posit that incorporating such real-time, meta-level supervision is a foundational component for building the next generation of truly scalable and reliable MAS.
7 Conclusion
@@ A.3.2 When to Supervise: The Prioritized Adaptive Filter
    Excessive Observation Length: Finally, if none of the above conditions are met, the filter checks if the length of the step.observations string exceeds a pre-defined threshold (3,000 characters in our implementation). If it does, the general type of Adaptive Observation Purification strategy is activated.
If none of these conditions are met, the step is approved by default, and no costly LLM-based supervision is invoked.
@@ 3. Action Execution
A.4 Case Study

### v2

@@ Abstract
While Multi-Agent Systems (MAS) excel at complex tasks, their growing autonomy with operational complexity often leads to critical inefficiencies, such as excessive token consumption and failures arising from misinformation. Existing methods primarily focus on post-hoc failure attribution, lacking proactive, real-time interventions to enhance robustness and efficiency. To this end, we introduce SupervisorAgent, a lightweight and modular framework for runtime, adaptive supervision that operates without altering the base agent’s architecture. Triggered by an LLM-free adaptive filter, SupervisorAgent intervenes at critical junctures to proactively correct errors, guide inefficient behaviors, and purify observations. On the challenging GAIA benchmark, SupervisorAgent reduces the token consumption of the Smolagent framework by an average of 29.68% without compromising its success rate. Extensive experiments across five additional benchmarks (math reasoning, code generation, and question answering) and various SoTA foundation models validate the broad applicability and robustness of our approach.
@@ 1 Introduction
The advent of powerful Large Language Models (LLMs) has catalyzed significant advancements in Multi-Agent Systems (MAS) (Liu et al., 2025a; Gao et al., 2025), enabling them to achieve remarkable performance across diverse and challenging domains such as mathematical reasoning (Shang et al., 2025), code generation (Lu et al., 2025), and complex question answering (Luo et al., 2025). This progress has spurred research into sophisticated agent architectures, including self-evolving systems that learn from feedback and experience (Shi et al., 2025b; Liu et al., 2025b), and dynamic topologies that adapt to task complexity (Li et al., 2025a; b). However, a critical paradox has emerged: as these systems grow more capable and complex, they often become less robust and economically viable (Wu et al., 2025a; Huang et al., 2025). Systemic inefficiencies incur prohibitive computational costs, while intricate interactions introduce vectors for unpredictable failures (Zhang et al., 2025e).
@@
Furthermore, the issue of economic inefficiency is a major barrier to the real-world deployment of MAS (Wang et al., 2025a). We identify two primary sources of this inefficiency. First, agents often struggle with long observations, such as verbose web pages or tool outputs, which flood their context windows. This not only inflates token costs but can also obscure critical information, causing the agent to lose focus and derail its task execution (Hosseini et al., 2025). Second, agents may adopt sub-optimal strategies, entering into repetitive action loops or choosing unnecessarily complex paths to a solution (Cemri et al., 2025), further wasting computational resources (see Figure 1a).
@@
    We validate the general applicability of our approach across five additional benchmarks spanning mathematical reasoning, code generation, and question answering. Our method consistently delivers substantial efficiency gains, highlighted by a 23.74% token reduction on HumanEval alongside an accuracy improvement. The framework’s effectiveness is further confirmed across various foundation models, including the GPT-4.1, Gemini-2.5-pro, and Qwen3 series.
@@ 4 Methodology
Building upon the formalism of a Supervised Multi-Agent System (SMAS) introduced in Section 3, we now detail the architecture and operational workflow of our SupervisorAgent (illustrated in Figure 2). Our methodology is structured around three fundamental questions: What to supervise, When to supervise, and How to supervise. We defer the specific implementation details, including all hyperparameters and prompts, to Appendix A.3 and A.7.
@@ 5.1 Experimental Setup
On several benchmarks, we compare SupervisorAgent against a comprehensive set of agentic systems equipped with web-browsing and code execution capabilities. These baselines fall into two categories: (1) single agent execution methods: including vanilla LLM, Self Consistency CoT (3 answers) (Wang et al., 2023), and CodeAgent (Roucher et al., 2025); and (2) multi-agent systems, including Smolagent (Roucher et al., 2025), OAgents (Zhu et al., 2025), MetaAgent (Zhang et al., 2025h), OWL (role playing) (Hu et al., 2025), and AWorld (Xie et al., 2025; Yu et al., 2025). Detailed descriptions of these baselines are provided in Appendix A.2.2.
@@
To assess model-agnosticism, we test SupervisorAgent with multiple foundation models. For the demanding GAIA benchmark, we primarily use GPT-4.1 as the base model for all agents, and evaluate SupervisorAgent when powered by GPT-4.1 (OpenAI, 2025), Gemini-2.5-pro-0605 (Team, 2025a), and Qwen3-235B-2507 (Team, 2025b). For all other benchmarks, we employ the efficient and powerful Qwen3-32B (Team, 2025b) for both the base agents and the SupervisorAgent to assess performance in a more resource-constrained setting.
@@
Table 1: Overall performance on the GAIA validation set. Our SMAS consistently reduces the average token cost comparing to Smolagent baseline while achieving competitive pass@k success rates.
@@ 5.2 Results and Analysis
Table 2: Generalization across diverse benchmarks. SupervisorAgent consistently reduces token costs while maintaining or improving accuracy on tasks spanning mathematical reasoning, code generation, and question answering. All reported gains are relative to the Smolagent baseline.
@@
We conducted an ablation study on the full GAIA validation set to isolate the impact of SupervisorAgent’s three core strategies (Table 3, Figure 4(a)). A comparison of the full framework with w/o Correction (Proactive Error Correction), w/o Guidance (Guidance for Inefficiency), and w/o Purification (Adaptive Observation Purification) reveals distinct roles. Purification is the primary driver of efficiency; disabling it drastically reduces token savings (from 29.68% to 15.96%). Conversely, removing Correction or Guidance results in the most significant accuracy drops, confirming their necessity for robustness. This underscores a synergistic design: while Purification minimizes cost, Correction and Guidance ensure task success, justifying their marginal overhead. These benefits are particularly pronounced on high-cost tasks (see Appendix A.4.2).
Table 3: Ablation study of SupervisorAgent’s components on the full GAIA validation set.
@@
Smolagent 50.91 527,759 298,506 619,591 691,331
+ SMAS (w/o Correction) 47.88 354,226 ↓32.88% 221,515 ↓25.79% 363,871 ↓41.27% 592,852 ↓14.24%
+ SMAS (w/o Guidance) 48.48 363,644 ↓31.10% 253,591 ↓15.05% 419,913 ↓32.23% 401,861 ↓41.87%
+ SMAS (w/o Purification) 49.70 443,520 ↓15.96% 270,058 ↓9.53% 502,937 ↓18.83% 600,582 ↓13.13%
+ SMAS 50.91 371,119 ↓29.68% 258,279 ↓13.48% 404,955 ↓34.64% 489,222 ↓29.23%
@@
To verify MAS-agnostic feature of SupervisorAgent, we integrate SupervisorAgent into two distinct multi-agent system frameworks: AWorld (Xie et al., 2025) and OAgents (Zhu et al., 2025), and evaluate their performance on the subset of GAIA benchmark(top-10 most token-intensive tasks per GAIA level). The results, presented in Table 4, indicate that SupervisorAgent consistently enhances the performance of both frameworks, underscoring its versatility and effectiveness across different MAS architectures.
Specifically, integrated with AWorld (Xie et al., 2025), our SMAS(AWorld) achieved superior average accuracy over both the original AWorld (without Guard) and the Guard-enabled version. Furthermore, SMAS(AWorld) demonstrated substantial token efficiency, saving 36.54% on average versus AWorld (with Guard). With savings reaching 48.38% on Level 3 tasks, it confirms SMAS’s ability to enhance tool-intensive MAS. Applying SupervisorAgent to OAgents (Zhu et al., 2025) further validated its general applicability, reducing average token consumption by 39.36% while maintaining competitive accuracy. Interestingly, the largest token reduction (50.19%) occurred on Level 1 tasks, exceeding the savings on Level 3 (40.63%). While this might reflect OAgents’ inherent proficiency on harder tasks, the significant overall savings underscore SupervisorAgent’s broad utility.
Overhead analysis.
Crucially, all efficiency gains reported in this work represent net savings, fully accounting for the cost of SupervisorAgent. As detailed in Table 6, the supervisor itself incurs a modest overhead, averaging only 15.45% of total token usage, which validates its lightweight design. Regarding latency, the supervisory interventions introduce an average increase of less than one minute and a half per task (Table 7). We consider this temporal cost a justifiable trade-off given the substantial economic savings achieved in complex multi-agent workflows. A comprehensive overhead analysis is provided in Appendix A.4.1.
Table 4: Cross-framework performance of SupervisorAgent. Evaluated on GAIA subset (top-10 most token-intensive tasks per level).
@@ 6 Discussion and Conclusion
These insights inform several promising avenues for future work. First, moving beyond heuristic rules, exploring a learning-based adaptive filter could enable more precise, dynamic control over supervisor invocations. This aligns with the broader goal of developing a self-evolving, memory-augmented version of SupervisorAgent. Second, further research should focus on mitigating the latency overhead introduced by supervisory calls to enhance real-time applicability, alongside creating sophisticated purification techniques that address the “noise-as-signal” trade-off. Finally, developing a universal resource consumption metric for MAS remains a critical open challenge. Ultimately, we posit that incorporating such real-time, meta-level supervision is a foundational component for building the next generation of truly scalable and reliable MAS.
Conclusion.
@@
Acknowledgements
This work was supported in part by the National Key Research and Development Program of China (2024YFF0907803), Research Fund for International Scientists of National Natural Science Foundation of China (72350710798), National Natural Science Foundation of China (NSFC) under No. 62576285, 62276230, Research Center for Industries of the Future (RCIF) at Westlake University, and Westlake Education Foundation.
Ethics Statement
Our work aims to improve the reliability and efficiency of Multi-Agent Systems, a crucial step for developing practical and beneficial autonomous technologies. We believe that by introducing a mechanism for real-time supervision, our framework provides a paradigm not only for performance optimization but also for enhancing the safety and predictability of future agentic systems. Our research was conducted on publicly available benchmarks, did not involve private user data, and adheres to the ICLR Code of Ethics.
Reproducibility Statement
We are committed to ensuring our work is reproducible. The core architecture and logic of SupervisorAgent are detailed in Section 4, with theoretical formalisms in Section 3. For direct replication, we provide all implementation details and final prompts in Appendix A.3A.7, and our code is available at https://github.com/LINs-lab/SupervisorAgent. The datasets and metrics used in our extensive experiments (Section 5) are all based on publicly available benchmarks, allowing for direct comparison and validation of our results.
@@ A.3 Implementation details
Table 5: Hyperparameter settings for the Heuristic-Based Adaptive Filter across different benchmarks. The symbols correspond to the definitions in Algorithm 1.
Condition Parameter (Symbol) GAIA HumanEval MBPP AIME DROP GSM-Hard
Inefficient Step Check Interval ($\tau_{\text{step}}$) 8 6 4 4 4 4
Loop Detection Window ($\tau_{\text{loop}}$) 5 5 3 3 3 3
Excessive Length Threshold ($\tau_{\text{len}}$) 3000
@@ A.3.2 When to Supervise: The Prioritized Adaptive Filter
    Excessive Observation Length: Finally, if none of the above conditions are met, the filter checks if the length of the step.observations string exceeds a pre-defined threshold $\tau_{\text{len}}$ (3,000 characters in our implementation). If it does, the general type of Adaptive Observation Purification strategy is activated.
If none of these trigger conditions are met, the step is approved by default, thereby avoiding any unnecessary LLM-based supervision overhead. The filter’s sensitivity is governed by three key hyperparameters: $\tau_{\text{step}}$ and $\tau_{\text{loop}}$ modulate the detection of inefficient behaviors, while $\tau_{\text{len}}$ defines the threshold for identifying excessive observations. The complete logic of this heuristic-based mechanism is formalized in Algorithm 1.
Algorithm 1 Heuristic-Based Adaptive Filter Logic
1:Current execution history $\mathcal{H}$ (sequence of steps), Current observation $o$, Error status $e$
2:Hyperparameters: $\tau_{\text{step}}$ (step check interval), $\tau_{\text{loop}}$ (loop detection window), $\tau_{\text{len}}$ (length threshold)
3:Boolean flag triggertrigger, Intervention context $c$
[…]
7:if $e$ is True then
8:  return $(\text{True},c_{\text{error}})$
9:end if
[…]
12:if $N>0$ and $N\pmod{\tau_{\text{step}}}=0$ then $\triangleright$ Periodic strategy check
13:  return $(\text{True},c_{\text{inefficient}})$
14:end if
15:if $N\geq\tau_{\text{loop}}$ then $\triangleright$ Repetitive loop detection
16:  $\mathcal{A}_{\text{recent}}\leftarrow\text{GetLastToolCalls}(\mathcal{H},\text{window}=\tau_{\text{loop}})$
17:  if $|\text{Unique}(\mathcal{A}_{\text{recent}})|=1$ then
18:   return $(\text{True},c_{\text{inefficient}})$
19:  end if
20:end if
[…]
22:if $\text{Length}(o)>\tau_{\text{len}}$ then
23:  return $(\text{True},c_{\text{excessive}})$
24:end if
25:return ((trigger, $c)$
Figure 5: Sensitivity analysis on excessive threshold ($\tau_{\text{len}}$). Evaluated on GAIA subset (top-10 most token-intensive tasks per level).
Sensitivity analysis and hyperparameter configuration.
Table 5 details the hyperparameter settings for our Heuristic-Based Adaptive Filter across different benchmarks. We conducted a sensitivity analysis on the excessive observation length threshold ($\tau_{\text{len}}$) using the representative subset of the GAIA benchmark (top-10 most token-intensive tasks per level), as illustrated in Figure 5. The results indicate that our method maintains robust performance across a wide range of threshold values, demonstrating its adaptability without significant degradation. Although slight performance peaks are observed at $\tau_{\text{len}}=1000$ and $\tau_{\text{len}}=5000$, we selected $3000$ as the default value. This choice strikes a prudent balance between sensitivity (catching enough noise) and specificity (preserving useful context), ensuring optimal performance across diverse tasks while minimizing the risk of over-intervention.
Configuration for OAgents.
For the OAgents framework (Zhu et al., 2025), we adjusted the parameters to $\tau_{\text{step}}=6$, $\tau_{\text{loop}}=3$, and $\tau_{\text{len}}=10000$. The significantly higher $\tau_{\text{len}}$ (compared to 3000 in Smolagent) is necessitated by OAgents’ architecture, which tends to generate extensive verbose outputs due to its complex tool usage and memory retrieval modules. A higher threshold is essential here to effectively identify truly excessive information without triggering false positives on standard OAgents operations.
Adaptation for AWorld: An MCP-Based Approach.
Unlike the external heuristic filter used in Smolagent, the integration with AWorld (Yu et al., 2025) leverages its native tool-use architecture. We implemented the SupervisorAgent as a Model Context Protocol (MCP) service, allowing it to be dynamically discovered and invoked by the AWorld agent. This adaptation involves three key modifications:
  • Mandatory Invocation via System Prompt: We refined AWorld’s system prompt to enforce a protocol where the agent must invoke the SupervisorAgent during critical phases—specifically “Information Gathering” and “Thinking Process Reviewing”. This ensures the SupervisorAgent acts as a mandatory gatekeeper for logical consistency.
  • Capability-Based Routing: We defined a specific MCP schema where the SupervisorAgent broadcasts its capabilities, including Error Root Cause Diagnosis, Structured Information Synthesis, and Workflow Efficiency Assessment. This allows the AWorld agent to match its current execution status (e.g., encountering an exception or synthesizing search results) with the appropriate Supervisor function.
  • Trigger Mechanism: Instead of counting token length, the trigger is semantic. Explicit error returns from AWorld’s system serve as triggers for error_analysis, while the sub_agent_result_synthesis and inefficiency_analysis modes are seamlessly integrated into the MCP process flow to facilitate output verification.
@@ 3. Action Execution
A.4 Extended Experimental Analysis
A.4.1 Overhead analysis of SupervisorAgent
Token overhead analysis
The token overhead of SupervisorAgent itself is shown in Table 6 and Figure 6. We analyze the token consumption on the GAIA validation set under different pass@k settings. The results indicate that integrating SupervisorAgent leads to a significant reduction in overall token usage across all complexity levels (L1, L2, L3) and pass@k configurations. Specifically, SupervisorAgent achieves an average token saving of 35.95% across all settings compared to the baseline Smolagent. This substantial decrease in token consumption highlights SupervisorAgent’s effectiveness in optimizing the multi-agent system’s efficiency by reducing unnecessary interactions and streamlining the reasoning process.
Notably, even when accounting for the additional tokens introduced by SupervisorAgent’s supervisory interventions, the net token consumption remains significantly lower than that of the baseline (already reported in main content). And the token overhead of SupervisorAgent only contains about 15.45% in average of the total tokens used in the Smolagent baseline. This demonstrates that the benefits of improved efficiency and reduced redundancy far outweigh the costs associated with supervision.
Table 6: Token efficiency analysis on GAIA validation set. Comparison of token consumption across different pass@k settings.
Method Avg. Tokens (K) L1 Tokens (K) L2 Tokens (K) L3 Tokens (K)
pass@1
Smolagent 527.76 298.51 619.59 691.33
+ Supervised MAS 314.07 ↓40.49% 220.63 ↓26.09% 342.18 ↓44.77% 411.58 ↓40.47%
+ Supervised MAS (NET) 371.12 ↓29.68% 258.28 ↓13.48% 404.96 ↓34.64% 489.22 ↓29.23%
pass@2
Smolagent 467.19 275.85 548.02 589.92
+ Supervised MAS 329.51 ↓29.47% 231.96 ↓15.91% 354.21 ↓35.37% 446.64 ↓24.29%
+ Supervised MAS (NET) 389.55 ↓16.62% 270.07 ↓2.10% 420.97 ↓23.18% 529.20 ↓10.29%
pass@3
Smolagent 502.40 282.14 605.05 611.87
+ Supervised MAS 312.06 ↓37.89% 236.28 ↓16.25% 342.36 ↓43.42% 366.31 ↓40.13%
+ Supervised MAS (NET) 369.52 ↓26.45% 276.84 ↓1.88% 409.05 ↓32.39% 427.72 ↓30.10%
Figure 6: SupervisorAgent overhead on the GAIA benchmark.
Latency Overhead Analysis.
Table 7 and Figure 7 analyzes the temporal impact of SupervisorAgent. While integrating the supervisor introduces an average latency increase of 37.27%, this translates to an absolute delay of less than 1.5 minutes per task. Crucially, the ablation study reveals that Adaptive Observation Purification is the primary driver of this latency. Notably, the w/o Purification variant exhibits a runtime nearly identical to the baseline (236.21s vs. 233.96s). This indicates that the overhead is strictly tied to the processing of excessive information. We consider this a strategic trade-off: exchanging a modest temporal cost for substantial economic (token) savings and enhanced system robustness.
Table 7: Ablation study of SupervisorAgent’s components regarding Latency on the GAIA validation set. Average latency (in seconds) is reported for different complexity levels.
Method Avg. Acc. Avg. Latency (s) Level 1 Avg. Lat. (s) Level 2 Avg. Lat. (s) Level 3 Avg. Lat. (s)
Smolagent 50.91 233.96 155.83 247.47 348.58
+ SMAS (w/o Correction) 47.88 280.77 ↑20.01% 193.98 ↑24.48% 276.50 ↑11.73% 471.81 ↑35.35%
+ SMAS (w/o Guidance) 48.48 271.95 ↑16.24% 211.04 ↑35.43% 291.22 ↑17.68% 332.38 ↓4.65%
+ SMAS (w/o Purification) 49.70 236.21 ↑0.96% 137.00 ↓12.08% 252.01 ↑1.83% 386.19 ↑10.79%
+ SMAS 50.91 321.15 ↑37.27% 233.11 ↑49.59% 320.93 ↑29.68% 501.31 ↑43.81%
Figure 7: SupervisorAgent latency on the GAIA benchmark.
A.4.2 Performance Analysis on Token-Intensive Scenarios
Complementing the comprehensive ablation study on the full GAIA benchmark (Table 3), this section zooms in on the most demanding scenarios: the top-10 most token-intensive tasks per GAIA level. This analysis aims to evaluate the scalability and robustness of SupervisorAgent under extreme computational loads.
Amplified Efficiency and Robustness.
As detailed in Table 8, the benefits of SupervisorAgent are significantly amplified in high-complexity regimes. While the average token reduction on the full dataset is 29.68%, SMAS achieves a remarkable 50.13% reduction on this intensive subset. More importantly, unlike the full dataset where accuracy remains stable, SMAS yields a distinct accuracy improvement of 6.67% (from 40.00% to 46.67%) on these hard tasks. This suggests that as task complexity and context length increase, the Supervisor’s interventions become indispensable not just for cost-saving, but for enabling the agent to complete tasks that were previously intractable due to context overflow or reasoning derailment.
Component Contribution in Extremes.
The ablation results on this subset reinforce the synergistic roles of our three strategies. The w/o Purification variant shows a drastic drop in efficiency (savings drop from 50.13% to roughly 40%), confirming that Adaptive Observation Purification is the primary countermeasure against the exponential token growth in complex tasks. Meanwhile, the removal of Correction or Guidance leads to a sharp decline in accuracy back to the baseline level (40.00%), verifying that these modules are the key safety guardrails that allow the system to navigate long-horizon tasks successfully.
Table 8: Ablation study of SupervisorAgent’s components on the subset of GAIA benchmark (top-10 most token-intensive tasks per GAIA level).
Smolagent 40.00 1,446,526 933,013 2,037,437 1,369,131
A.5 Failure Mode Analysis
While SupervisorAgent demonstrated robustness across benchmarks, it relies on backbone LLMs and is thus subject to their inherent limitations. We identify three primary failure modes and their implications.
Information Loss during Purification.
The Adaptive Observation Purification module faces a trade-off between context reduction and information preservation. In extreme cases observed in the OAgents framework, where single observations exceeded 200,000 characters, the Supervisor risks hallucinating or omitting critical details during compression. However, our empirical results suggest that the system’s resilience to context overflow generally outweighs the cost of granular information loss, as evidenced by the overall token savings and success rates (see in Table 4).
Ineffective Guidance and Loops.
The Inefficiency Guidance module may occasionally provide suboptimal advice or fail to break a stubborn loop. To mitigate the risk of the Supervisor itself becoming a source of latency (e.g., engaging in an infinite correction loop with a non-responsive agent), we enforce a hard constraints of maximum two guidance interventions per sub-task. While this design prioritizes bounded latency over guaranteed resolution for the hardest tasks, it effectively prevents runaway costs.
Variance in Trigger Frequency across Backbones.
Contrary to the assumption that a supervisor acts uniformly, our analysis reveals that SupervisorAgent exhibits different operational behaviors depending on the backbone model’s capability. For instance, as shown in our logs with Qwen3-235B, less capable models trigger the Error Correction module significantly more often due to frequent basic failures (e.g., malformed JSON tool calls). This frequent firing increases the Supervisor’s token overhead, partially explaining the lower net token savings compared to GPT-4.1 (see in Figure 4(b)). Conversely, stronger models may trigger supervision too rarely if the heuristic filter is not sensitive enough to their subtle logic errors. This highlights that while the framework is model-agnostic, the efficiency gains are correlated with the backbone LLM’s adherence to instruction following.
A.6 Case Study